# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v2

@@ 3 Pólya’s conjecture on a class of domains
Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture, $n\geq 2$, $\Omega_{2}$ is an interval in ${\mathbb{R}}$ and $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega_{1}\times\Omega_{2}$. There exists $C_{P}=C(\Omega_{3},n)$ such that if $|\Omega_{1}|\geq C_{P}$ then the Pólya conjecture holds on $\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$.
@@
(3.1) $\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$
@@
Note that (3.1) can be stated for all $\lambda>0$ without referring to $\Lambda$, but generally with larger constant $C(\Omega_{3})$. We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large constant using the subtracting argument, will give better estimate, see Remark 3.5 below.
@@ 3.1 The case of two dimension
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
@@
Using (2.3) one can give a lower bound for $\mathcal{N}^{D}_{\mathcal{R}}(\lambda)$ for any rectangle $\mathcal{R}$ in ${\mathbb{R}}^{2}$. Inductively, a lower bound for the eigenvalue counting function can be explicitly given on any cuboid in higher dimension.
@@
Denote by $\Omega_{0}=\Omega_{1}\times\Omega_{2}$ and $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\cup_{k=1}^{\infty}Q_{k}}$. Let
@@
Then
@@
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{2}}{3|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
@@
$k_{0}=\left\lfloor\frac{|\Omega|}{4\pi}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{2}{3|\Omega_{2}|}\right]^{2}\right\rfloor=\left\lfloor\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}\right\rfloor,$
the first $k_{0}$ Dirichlet eigenvalues of $\Omega$ satisfy Pólya’s conjecture, that is, for $1\leq i\leq k_{0}$,
$\frac{|\Omega|\omega(2)}{(2\pi)^{2}}\lambda_{i}(\Omega)\geq i.$
In particular, since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}\geq 2\sqrt{2}\pi\mathscr{V}_{Q}$,
$|\Omega|=|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1)\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1),$
@@
$\lambda_{k_{0}}(\Omega)\geq\frac{4\pi}{|\Omega|}\left(\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}-1\right)>33.$
Recall that Lemma 2.1 implies that for $\lambda>\pi^{2}/|\Omega_{1}|^{2}+\pi^{2}/|\Omega_{2}|^{2}$,
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}$
$\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right).$
In particular, for $\lambda>33$, it holds that
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\frac{3}{2}\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}}\right)$
(3.2) $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{1}{|\Omega_{2}|\sqrt{\lambda}}\right).$
@@
$|\Omega_{1}\times\Omega_{2}|=|\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}|.$
So by monotonicity of Dirichlet eigenvalues, Pólya’s conjecture holds on $\Omega\cup\cup_{1\leq k<\infty}Q_{k}$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda).$
Moreover, by Lemma 3.3 (ii), we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
@@
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{1}}{|\Omega_{2}|\sqrt{\lambda}}\right)-\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right)$
$=\frac{|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right)$
$=\frac{|\Omega|}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right).$
Since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}$, we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda,$
@@ 3.2 Three and higher dimensions
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}+c^{2}),\,a,b,c\in\mathbb{N}\right\}.$
@@
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}\sum_{i=1}^{n}a_{i}^{2},\,a_{i}\in\mathbb{N}\right\}.$
@@
(3.3) $C_{1}(n-1)=\frac{3\pi\mathscr{B}(\frac{n-1}{2}-1,2)}{2\mathscr{B}(\frac{n-1}{2}-1,5/2)}C_{3}(\frac{n-1}{2})=\frac{3\pi\mathscr{B}(\frac{n-3}{2},2)}{2\mathscr{B}(\frac{n-3}{2},5/2)}C_{3}(\frac{n-1}{2}),$
@@
Let $n\geq 3$ and suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture (e.g. ball, tiling domain, or our example in ${\mathbb{R}}^{n-1}$), $\Omega_{2}$ is an interval in ${\mathbb{R}}$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$ that is contained in $\Omega_{1}\times\Omega_{2}$. If $|\Omega_{1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $(\Omega_{1}\times\Omega_{2})\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where $C_{2}(n-1)$ is as in (1.12),
@@
Recall that Lemma 2.1 implies that for $\lambda>\lambda_{1}(\Omega_{2})$,
@@
(3.4) $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$
@@
(3.5) $\mathcal{N}^{D}_{\Omega_{0}}(\lambda)=\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right).$
@@
(3.6) $=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}},$
@@ 3.3 Some other class satisfying Pólya’s conjecture
(3.7) $k\leq\begin{cases}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},\,&\,k\leq 2,\\ \frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},\,&\,k>2.\end{cases}$
@@
(3.8) $k\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda_{k}^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}^{\frac{n-1}{2}},$
@@
(3.9) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{0})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{0}|^{2/n}\omega(n)^{2/n}}+\frac{\pi}{|\Omega_{2}|^{2}},$
@@
4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate

### v3

@@ 3 Pólya’s conjecture on a class of domains
(i) Suppose that $\Omega$ is a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Suppose that $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega$. There exists $C_{P}=C(\Omega_{3},n)$ such that if $|S_{n}(\Omega)|\geq C_{P}$ then the Pólya conjecture holds on $\Omega\setminus\overline{\Omega_{3}}$.
(ii) Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture, $n\geq 2$, $\Omega_{2}$ is an interval in $\mathbb{R}$} and $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega_{1}\times\Omega_{2}$. There exists $C_{P}=C(\Omega_{3},n)$ such that if $|\Omega_{1}|\geq C_{P}$ then the Pólya conjecture holds on $\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$.
@@
We only prove (ii), the proof of (i) is the same by using Theorem 1.10.
@@
$\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$ (3.1)
@@
Note that (3.1) can be stated for all $\lambda>0$ without referring to $\Lambda$, but generally with larger constant $C(\Omega_{3})$. We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large eigenvalues using the subtracting argument, will give better estimate, see Remark 3.5 below.
@@ 3.1 The case of two dimension
$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
@@
Using (2.3) one can give a lower bound for $\mathcal{N}^{D}_{\mathscr{R}}(\lambda)$ for any rectangle $\mathscr{R}$ in ${\mathbb{R}}^{2}$. Inductively, a lower bound for the eigenvalue counting function can be explicitly given on any cuboid in higher dimension.
@@
Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}\times(0,w(\Omega))$, and let $S_{n}(\Omega)$ be the surface of a side face of $\Omega$ in the $n$-th direction. Denote by $\Omega_{0}=\Omega\setminus\overline{\cup_{k=1}^{\infty}Q_{k}}$. Let
@@
Then by the definition of $\{Q_{k}\}$, it holds that $w(\Omega)\geq 1$ and
@@
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{{2}}{3|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
@@
$k_{0}=\left\lfloor\frac{|\Omega_{0}|}{4\pi}\left[\left(1-\frac{|\Omega_{0}|}{|\Omega|}\right)^{-1}\frac{2}{3|w(\Omega)|}\right]^{2}\right\rfloor=\left\lfloor\frac{|\Omega_{0}|}{4\pi}\left[\frac{2|S_{n}(\Omega)|}{3\mathscr{V}_{Q}}\right]^{2}\right\rfloor,$
the first $k_{0}$ Dirichlet eigenvalues of $\Omega_{0}$ satisfy Pólya’s conjecture, that is, for $1\leq i\leq k_{0}$,
$\frac{|\Omega_{0}|\omega(2)}{(2\pi)^{2}}\lambda_{i}(\Omega)\geq i.$
In particular, since $|S_{n}(\Omega)|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}\geq 2\sqrt{2}\pi\mathscr{V}_{Q}$,
$|\Omega_{0}|=|\Omega|-\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1)\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1),$
@@
$\lambda_{k_{0}}(\Omega)\geq\frac{4\pi}{|\Omega|}\left(\frac{|\Omega|}{4\pi}\left[\frac{2|S_{n}(\Omega)|}{3\mathscr{V}_{Q}}\right]^{2}-1\right)>33.$
By Remark 2.5, we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{2}{|w(\Omega)|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|w(\Omega)|^{2}\lambda}\sqrt{\lfloor\frac{|w(\Omega)|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}.$
This together with $|w(\Omega)|\geq 1$ implies that for $\lambda>33$, it holds
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{1}{|w(\Omega)|\sqrt{\lambda}}\right).$ (3.2)
@@
$|\Omega|=|\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}|.$
So by monotonicity of Dirichlet eigenvalues, Pólya’s conjecture holds on $\Omega_{0}\cup\cup_{1\leq k<\infty}Q_{k}$, and
$\mathcal{N}^{D}_{\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega}(\lambda).$
Moreover, by Lemma 3.3 (ii), we see that for $\lambda>0$,
$\mathcal{N}^{D}_{\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega_{0}}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega_{0}}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
@@
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{1}{|w(\Omega)|\sqrt{\lambda}}\right)-\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right)$
$=\frac{|\Omega|-\mathscr{V}_{Q}}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|S_{n}(\Omega)|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right)$
$=\frac{|\Omega_{0}|}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|S_{n}(\Omega)|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right).$
Since $|S_{n}(\Omega)|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}$, we see that
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{0}|}{4\pi}\lambda,$
@@ 3.2 Three and higher dimensions
$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}+c^{2}),\,a,b,c\in\mathbb{N}\right\}.$
@@
$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}\sum_{i=1}^{n}a_{i}^{2},\,a_{i}\in\mathbb{N}\right\}.$
@@
$C_{1}(n-1)=\frac{3\pi\mathscr{B}(\frac{n-1}{2}-1,2)}{2\mathscr{B}(\frac{n-1}{2}-1,5/2)}C_{3}(\frac{n-1}{2})=\frac{3\pi\mathscr{B}(\frac{n-3}{2},2)}{2\mathscr{B}(\frac{n-3}{2},5/2)}C_{3}(\frac{n-1}{2}),$ (3.3)
@@
Let $n\geq 3$.
(i) Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture (e.g. ball, tiling domain, or our example in ${\mathbb{R}}^{n-1}$), $\Omega_{2}$ is an interval in $\mathbb{R}$}. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$} that is contained in $\Omega_{1}\times\Omega_{2}$. If $|\Omega_{1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $(\Omega_{1}\times\Omega_{2})\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where $C_{2}(n-1)$ is as in (1.12),
@@
(ii) Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$. Let $S_{n}(\Omega)$ be the surface of a side face of $\Omega$ in the $n$-th direction. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$} that is contained in $\Omega$. If $|S_{n}(\Omega)|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$,
@@
We only prove (i), the proof of (ii) is the same by using Theorem 1.10.
@@
Recall that Remark 2.5 implies that for $\lambda>\lambda_{1}(\Omega_{2})$,
@@
$\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$ (3.4)
@@
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)=\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right).$ (3.5)
@@
$=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}},$ (3.6)
@@ 3.3 Some other class satisfying Pólya’s conjecture
$k\leq\begin{cases}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},\,&\,k\leq 2,\\ \frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},\,&\,k>2.\end{cases}$ (3.7)
@@
$k\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda_{k}^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}^{\frac{n-1}{2}},$ (3.8)
@@
$\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{0})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{0}|^{2/n}\omega(n)^{2/n}}+\frac{\pi}{|\Omega_{2}|^{2}},$ (3.9)
@@
4 Quantitative remainder estimate